OpenStudy (anonymous):

Please help!!! =( Elimination method 3x + 2y= 9 6x + 5y= 15

OpenStudy (anonymous):

please tell me why i have to multiply it by 2

OpenStudy (kropot72):

If you multiply both sides of the first equation by 2, the coefficient of x will become the same as the coefficient of x in the second equation. Then, by subtracting the equations, the term in x will be eliminated and you can solve to find the value of y.

OpenStudy (anonymous):

i really dont understand y it has to be multiplied by 2

OpenStudy (anonymous):

where does the 2 come from and can it be a number other than 2

OpenStudy (kropot72):

We want the term in x in the first equation to become equal to the term in x in the second equation. So, what do we need to multiply 3x by to make it equal to 6x?

OpenStudy (anonymous):

how would it be equal to the second equation.......

OpenStudy (anonymous):

4y does not equal to 5y

OpenStudy (kropot72):

No, we are not trying to make the first equation equal to the second equation. We want to make the terms in x the same in both equations.

OpenStudy (anonymous):

im still confused...

OpenStudy (kropot72):

Can I repeat my earlier question: We want the term in x in the first equation to become equal to the term in x in the second equation. So, what do we need to multiply 3x by to make it equal to 6x?

OpenStudy (anonymous):

2

OpenStudy (kropot72):

Correct! So now we need to multiply all the terms in the first equation by 2, giving: 6x + 4y = 18 ...........(1) Now if we subtract the second equation from equation (1) we get: 6x + 4y -(6x + 5y) = 18 - 15 ........(2) Now can you simplify equation (2) to form a new equation with the terms in x eliminated?

OpenStudy (anonymous):

okay i still dont understand why it has to be 2..........

OpenStudy (kropot72):

We have chosen 2 as the multiplier for the first equation to get the result in equation (2) above, where the terms in x will be eliminated.

OpenStudy (kropot72):

If we had chosen a different multiplier, say 5, then we would not have been able to eliminate the terms in x simply by subtracting one equation from another.